OpenStudy (anonymous):

Evaluate the expression if t = 12, w = 10, x = 9, y = 4, and z = 3 tw divided by z+ xy=

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

u need to substitue the values and evaluate... to avoid confusion, u evaluate the numerator and denominator seeperately and then divide. so, NR= z+xy=3+9*4=39 DR=tw=12*10=120 so the ans=39/120=13/40.

OpenStudy (anonymous):

40/13

OpenStudy (anonymous):

its tw by z + xy

OpenStudy (anonymous):

its 40/13 in fraction it is 12*10/3+9*4 use BODMAS and do not cancel if it is addition

OpenStudy (anonymous):

yea sorry him is right . :)